# on the Radon-Nikodym Derivative

### v1

@@ IV Advanced folklore theorems
This section requires some additional notation. In particular, denote by $\triangle\left(\mathcal{X},\mathscr{F}_{\mathcal{X}}\right)$, or simply $\triangle\left(\mathcal{X}\right)$, the set of all probability measures on the measurable space $\left(\mathcal{X},\mathscr{F}_{\mathcal{X}}\right)$, where $\mathscr{F}_{\mathcal{X}}$ is a $\sigma$-algebra on $\mathcal{X}$. Using this notation, conditional probability measures can be defined as follows.
Definition 1 (Conditional Probability)
A family $P_{Y|X}~{}\triangleq~{}(P_{Y|X=x})_{x\in\mathcal{X}}$ of elements of $\triangle(\mathcal{Y},\mathscr{F}_{\mathcal{Y}})$ indexed by $\mathcal{X}$ is said to be a conditional probability measure if, for all sets $\mathcal{A}\in\mathscr{F}_{\mathcal{Y}}$, the map
$\mathcal{X}$$\to$ $[0,1]$ (35)
$x$ $\mapsto$ $P_{Y|X=x}(\mathcal{A})$ (36)
is Borel measurable. The set of such conditional probability measures is denoted by $\triangle(\mathcal{Y}|\mathcal{X})$.
A conditional probability $P_{Y|X}\in\triangle(\mathcal{Y}|\mathcal{X})$ and a probability measure $P_{X}\in\triangle(\mathcal{X})$ determine two unique probability measures in $\triangle(\mathcal{X}\times\mathcal{Y})$ and $\triangle(\mathcal{Y}\times\mathcal{X})$, respectively. These probability measures are denoted by $P_{XY}$ and $P_{YX}$, respectively, and for all sets $\mathcal{A}\in\mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}$, it follows that
$P_{XY}\left(\mathcal{A}\right)$ $=$ $\int P_{Y|X=x}\left(\mathcal{A}_{x}\right)\mathrm{d}P_{X}\left(x\right),$ (37)
@@
$\mathcal{A}_{x}\triangleq\left\{y\in\mathcal{Y}:(x,y)\in\mathcal{A}\right\}.$ (38)
@@
$P_{YX}\left(\mathcal{B}\right)$ $=$ $\int P_{Y|X=x}\left(\mathcal{B}_{x}\right)\mathrm{d}P_{X}\left(x\right),$ (39)
where $\mathcal{B}_{x}$ is the section of the set $\mathcal{B}$ determined by $x$. For all sets $\mathcal{A}\in\mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}$, let the set $\hat{\mathcal{A}}\in\mathscr{F}_{\mathcal{Y}}\times\mathscr{F}_{\mathcal{X}}$ be such that
$\hat{\mathcal{A}}$ $=$ $\left\{(y,x)\in\mathcal{Y}\times\mathcal{X}:(x,y)\in\mathcal{A}\right\}.$ (40)
@@
$P_{XY}\left(\mathcal{A}\right)$ $=$ $P_{YX}\left(\hat{\mathcal{A}}\right).$ (41)
Using this notation, the notion of marginal probability measures can be introduced as follows.
Definition 2 (Marginals)
Given two joint probability measures $P_{XY}\in\triangle\left(\mathcal{X}\times\mathcal{Y}\right)$ and $P_{YX}\in\triangle\left(\mathcal{Y}\times\mathcal{X}\right)$, satisfying (41), the marginal probability measures in $\triangle\left(\mathcal{X}\right)$ and $\triangle\left(\mathcal{Y}\right)$, denoted by $P_{X}$ and $P_{Y}$, respectively satisfy for all sets $\mathcal{A}\in\mathscr{F}_{\mathcal{X}}$ and for all sets $\mathcal{B}\in\mathscr{F}_{\mathcal{Y}}$,
$P_{X}\left(\mathcal{A}\right)$ $\triangleq P_{XY}\left(\mathcal{A}\times\mathcal{Y}\right)=$ $P_{YX}\left(\mathcal{Y}\times\mathcal{A}\right);\mbox{ and }$ (42)
$P_{Y}\left(\mathcal{B}\right)$ $\triangleq P_{XY}\left(\mathcal{X}\times\mathcal{B}\right)=$ $P_{YX}\left(\mathcal{B}\times\mathcal{X}\right).$ (43)
From the total probability theorem [5, Theorem 4.5.2], it follows that for all $\mathcal{A}\in\mathscr{F}_{\mathcal{Y}}$,
$P_{Y}(\mathcal{A})=\int\int_{\mathcal{A}}\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x);$ (44)
and for all $\mathcal{B}\in\mathscr{F}_{\mathcal{X}}$,
$P_{X}(\mathcal{B})=\int\int_{\mathcal{B}}\mathrm{d}P_{X|Y=y}(x)\mathrm{d}P_{Y}(y).$ (45)
The joint probability measures $P_{XY}$ and $P_{YX}$ can be described via the conditional probability measure $P_{Y|X}$ and the probability measure $P_{X}$ as in (37) and in (39); or via the conditional probability measure $P_{X|Y}\in\triangle\left(\mathcal{X}|\mathcal{Y}\right)$ and the marginal probability measure $P_{Y}\in\triangle\left(\mathcal{Y}\right)$. More specifically, for all sets $\mathcal{A}\in\mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}$, it follows that
@@
Theorem 9 (Unit Measure)
@@
$P_{Y}(\mathcal{A})$ $=$ $\int\int_{\mathcal{A}}\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x)$ (50)
$=$ $\int\int_{\mathcal{A}}\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}(y)\mathrm{d}P_{Y}(y)\mathrm{d}P_{X}(x)$ (51)
@@
The following folklore theorem is reminiscent of the Bayes rule.
Theorem 10 (Bayes-like rule)
Consider the conditional probability measures $P_{Y|X}$ and $P_{X|Y}$; the probability measures $P_{Y}$ and $P_{X}$ that satisfy (44) and (45); and the joint probability measures $P_{YX}$ and $P_{XY}$ in (39) and (46) respectively. Let also $P_{X}P_{Y}\in\triangle\left(\mathcal{X}\times\mathcal{Y},\mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}\right)$ and $P_{Y}P_{X}\in\triangle\left(\mathcal{Y}\times\mathcal{X},\mathscr{F}_{\mathcal{Y}}\times\mathscr{F}_{\mathcal{X}}\right)$ be the measures formed by the product of the marginals $P_{X}$ and $P_{Y}$. Assume that:
  • $\left(a\right)$
    For all $x\in\mathcal{X}$${P_{Y|X=x}}\ll{P_{Y}}$; and
  • $\left(b\right)$
    For all $y\in\mathcal{Y}$${P_{X|Y=y}}\ll{P_{X}}$.
Then,
@@
=$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{X}P_{Y}}$ $\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)$ (54)
=$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{X}P_{Y}}$ $\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right).$ (55)
@@
Note that assumptions $\left(a\right)$ and $\left(b\right)$ are sufficient for the Radon-Nikodym derivatives of $P_{XY}$ with respect to $P_{X}P_{Y}$ and $P_{YX}$ with respect to $P_{Y}P_{X}$ to exist. Hence, it follows that for all sets $\mathcal{A}~{}\in~{}\mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}$,
@@
$P_{XY}\left(\mathcal{A}\right)$ $=$ $\int\int_{\mathcal{A}_{y}}\mathrm{d}P_{X|Y=y}\left(x\right)\mathrm{d}P_{Y}\left(y\right)$ (57)
$=$ $\int\int_{\mathcal{A}_{y}}\frac{\mathrm{d}P_{X|Y=y}}{\mathrm{d}P_{X}}\left(x\right)\mathrm{d}P_{X}\left(x\right)\mathrm{d}P_{Y}\left(y\right)$ (58)
$=$ $\int\int\mathds{1}_{\mathcal{A}_{y}}(x)\frac{\mathrm{d}P_{X|Y=y}}{\mathrm{d}P_{X}}\left(x\right)\mathrm{d}P_{X}\left(x\right)\mathrm{d}P_{Y}\left(y\right)$ (59)
@@
$P_{XY}\left(\mathcal{A}\right)$ $=$ $\int\int_{\mathcal{A}_{x}}\mathrm{d}P_{Y|X=x}\left(y\right)\mathrm{d}P_{X}\left(x\right)$ (62)
$=$ $\int\int_{\mathcal{A}_{x}}\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)\mathrm{d}P_{Y}\left(y\right)\mathrm{d}P_{X}\left(x\right)$ (63)
$=$ $\int\int\mathds{1}_{\mathcal{A}_{x}}(y)\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)\mathrm{d}P_{Y}\left(y\right)\mathrm{d}P_{X}\left(x\right)$ (64)
$=$ $\int\int\mathds{1}_{\mathcal{A}_{y}}(x)\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)\mathrm{d}P_{X}\left(x\right)\mathrm{d}P_{Y}\left(y\right)$ (65)
@@
$P_{XY}\left(\mathcal{A}\right)=\int_{\hat{\mathcal{A}}}\mathrm{d}P_{YX}\left(y,x\right)$ (68)
$=$ $\int_{\hat{\mathcal{A}}}\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{Y}P_{X}\left(y,x\right)$
$=$ $\int\mathds{1}_{\hat{\mathcal{A}}}(y,x)\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{Y}P_{X}\left(y,x\right)$ (69)
$=$ $\int\int\mathds{1}_{\hat{\mathcal{A}}_{x}}(y)\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{Y}(y)\mathrm{d}P_{X}(x)$ (70)
$=$ $\int\int\mathds{1}_{\hat{\mathcal{A}}_{y}}(x)\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{X}(x)\mathrm{d}P_{Y}(y)$ (71)
$=$ $\int\mathds{1}_{\hat{\mathcal{A}}}(x,y)\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{X}P_{Y}(x,y)$ (72)
@@
where the set $\hat{\mathcal{A}}$ is defined in (40). By performing a change of measure using Theorem 2 and assumption $\left(a\right)$, the equality in (68) is obtained; and the equality in (71) follows by exchanging the order of integration [5, Theorem 2.6.6].
The proof is completed from Theorem 1 and by combining equations (56),  (61),  (67) and (73), which establish (53), (54) and (55). ∎
Theorem 11 (Inverse Bayes-like Rule)
Consider the conditional probability measures $P_{Y|X}$ and $P_{X|Y}$; and the probability measures $P_{Y}$ and $P_{X}$ that satisfy (44) and (45); and the joint probability measures $P_{YX}$ and $P_{XY}$ in (39) and (46) respectively. Assume that:
  • $\left(a\right)$
    For all $x\in\mathcal{X}$${P_{Y}}\ll{P_{Y|X=x}}$; and
  • $\left(b\right)$
    For all $y\in\mathcal{Y}$${P_{X}}\ll{P_{X|Y=y}}$.
Then,
$\frac{\mathrm{d}P_{X}P_{Y}}{\mathrm{d}P_{XY}}\left(x,y\right)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{XY}}\frac{\mathrm{d}P_{X}}{\mathrm{d}P_{X|Y=y}}\left(x\right)$ (74)
=$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{XY}}$ $\frac{\mathrm{d}P_{Y}}{\mathrm{d}P_{Y|X=x}}\left(y\right)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{YX}}\frac{\mathrm{d}P_{Y}P_{X}}{\mathrm{d}P_{YX}}\left(y,x\right).$
@@
V Conclusions and Final Remarks
This paper provides proofs for several well-known folklore theorems, as well as some other lesser-known results in information theory involving the Radon-Nikodym derivative. Notably, these theorems serve as fundamental tools for establishing various properties of Shannon’s information measures in a unified framework, that is, regardless of whether the underlying measures admit a probability mass function or a probability density function. Nonetheless, the proof of these properties are left out of the scope of this paper, due to space constraints. Finally, it is important to highlight that the presented results hold in full generality, as the reference measure is assumed to be any $\sigma$-finite measure.

### v2

@@ IV Advanced folklore theorems
This section focuses exclusively on probability measures with the aim of establishing closer ties to information measures, which requires some additional notation. In particular, denote by $\triangle\bigl{(}\mathcal{X},\mathscr{F}_{\mathcal{X}}\bigr{)}$, or simply $\triangle\bigl{(}\mathcal{X}\bigr{)}$, the set of all probability measures on the measurable space $\bigl{(}\mathcal{X},\mathscr{F}_{\mathcal{X}}\bigr{)}$, where $\mathscr{F}_{\mathcal{X}}$ is a $\sigma$-algebra on $\mathcal{X}$. Using this notation, conditional probability measures can be defined as follows.
Definition 1 (Conditional Probability).
A family $P_{Y|X}\,\triangleq\,(P_{Y|X=x})_{x\in\mathcal{X}}$ of elements of $\triangle\bigl{(}\mathcal{Y},\mathscr{F}_{\mathcal{Y}}\bigr{)}$ indexed by $\mathcal{X}$ is said to be a conditional probability measure if, for all sets $\mathcal{B}\in\mathscr{F}_{\mathcal{Y}}$, the map
[…]
is Borel measurable. The set of such conditional probability measures is denoted by $\triangle(\mathcal{Y}\!\mid\!\mathcal{X})$.
A conditional probability $P_{Y|X}\in\triangle(\mathcal{Y}\mid\mathcal{X})$ and a probability measure $P_{X}\in\triangle(\mathcal{X})$ jointly determine two unique probability measures in $\triangle(\mathcal{X}\times\mathcal{Y})$ and $\triangle(\mathcal{Y}\times\mathcal{X})$, respectively. These probability measures are denoted by $P_{XY}$ and $P_{YX}$, respectively, and for all sets $\mathcal{A}\in\mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}$, it follows that
$P_{XY}\bigl{(}\mathcal{A}\bigr{)}$ $=$ $\int P_{Y|X=x}\bigl{(}\mathcal{A}_{x}\bigr{)}\,\mathrm{d}P_{X}(x),$ (24)
@@
$\mathcal{A}_{x}$$\triangleq$ $\bigl{\{}y\in\mathcal{Y}:(x,y)\in\mathcal{A}\bigr{\}}.$ (25)
@@
$P_{YX}\bigl{(}\mathcal{B}\bigr{)}=\int P_{Y|X=x}\bigl{(}\mathcal{B}_{x}\bigr{)}\,\mathrm{d}P_{X}(x),$ (26)
where $\mathcal{B}_{x}$ is the section of the set $\mathcal{B}$ determined by $x$. Note that $P_{XY}$ is a measure in $\triangle\left(\mathcal{X}\times\mathcal{Y}\right)$ whereas $P_{YX}$ is a measure in $\triangle\left(\mathcal{Y}\times\mathcal{X}\right)$. Consequently, obtaining $P_{XY}$ from $P_{YX}$ requires more than simply renaming $X$ and $Y$; one must also properly redefine the corresponding measurable spaces. In particular, given a set $\mathcal{A}\in\mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}$, let the set $\widehat{\mathcal{A}}\in\mathscr{F}_{\mathcal{Y}}\times\mathscr{F}_{\mathcal{X}}$ be such that
$\widehat{\mathcal{A}}=\bigl{\{}(y,x)\in\mathcal{Y}\times\mathcal{X}:(x,y)\in\mathcal{A}\bigr{\}}.$ (27)
@@
$P_{XY}\bigl{(}\mathcal{A}\bigr{)}=P_{YX}\bigl{(}\widehat{\mathcal{A}}\bigr{)}.$ (28)
Using this notation, a marginal probability measure is defined as follows.
Definition 2 (Marginal).
Given two joint probability measures $P_{XY}\in\triangle\left(\mathcal{X}\times\mathcal{Y}\right)$ and $P_{YX}\in\triangle\left(\mathcal{Y}\times\mathcal{X}\right)$, satisfying (28), the marginal probability measures in $\triangle\left(\mathcal{X}\right)$ and $\triangle\left(\mathcal{Y}\right)$, denoted by $P_{X}$ and $P_{Y}$, respectively, satisfy for all sets $\mathcal{A}\in\mathscr{F}_{\mathcal{X}}$
$P_{X}\left(\mathcal{A}\right)\triangleq P_{XY}\left(\mathcal{A}\times\mathcal{Y}\right)=P_{YX}\left(\mathcal{Y}\times\mathcal{A}\right);$ (29)
and for all sets $\mathcal{B}\in\mathscr{F}_{\mathcal{Y}}$,
$P_{Y}\left(\mathcal{B}\right)\triangleq P_{YX}\left(\mathcal{B}\times\mathcal{X}\right)=P_{XY}\left(\mathcal{X}\times\mathcal{B}\right).$ (30)
From the total probability theorem [5, Theorem 4.5.2], it follows that for all $\mathcal{B}\in\mathscr{F}_{\mathcal{Y}}$,
$P_{Y}(\mathcal{B})=\iint_{\mathcal{B}}\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x);$ (31)
and for all $\mathcal{A}\in\mathscr{F}_{\mathcal{X}}$,
$P_{X}(\mathcal{A})=\iint_{\mathcal{A}}\mathrm{d}P_{X|Y=y}(x)\mathrm{d}P_{Y}(y).$ (32)
In a nutshell, the joint probability measures $P_{XY}$ and $P_{YX}$ can be described via the conditional probability measure $P_{Y|X}$ and the probability measure $P_{X}$ as in (24) and in (26); or via the conditional probability measure $P_{X|Y}\in\triangle\left(\mathcal{X}|\mathcal{Y}\right)$ and the marginal probability measure $P_{Y}\in\triangle\left(\mathcal{Y}\right)$. More specifically, for all sets $\mathcal{A}\in\mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}$, it follows that
@@
Theorem 9 (Unit Measure).
@@
$P_{Y}(\mathcal{A})$ $=$ $\iint_{\mathcal{A}}\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x)$ (37)
$=$ $\iint_{\mathcal{A}}\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}(y)\mathrm{d}P_{Y}(y)\mathrm{d}P_{X}(x)$ (38)
@@
The following folklore theorems are reminiscent of the Bayes rule.
Theorem 10 (Bayes-like rule).
Consider the conditional probability measures $P_{Y|X}$ and $P_{X|Y}$; the probability measures $P_{Y}$ and $P_{X}$ that satisfy (31) and (32); and the joint probability measures $P_{YX}$ and $P_{XY}$ in (26) and (33) respectively. Let also $P_{X}P_{Y}\in\triangle\left(\mathcal{X}\times\mathcal{Y},\mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}\right)$ and $P_{Y}P_{X}\in\triangle\left(\mathcal{Y}\times\mathcal{X},\mathscr{F}_{\mathcal{Y}}\times\mathscr{F}_{\mathcal{X}}\right)$ be the measures formed by the product of the marginals $P_{X}$ and $P_{Y}$. Assume that: $\left(a\right)$ For all $x\in\mathcal{X}$${P_{Y|X=x}}\ll{P_{Y}}$; and $\left(b\right)$ For all $y\in\mathcal{Y}$${P_{X|Y=y}}\ll{P_{X}}$. Then,
@@
=$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{X}P_{Y}}\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)$ =$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{X}P_{Y}}$ $\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right).$ (41)
@@
Note that assumptions $\left(a\right)$ and $\left(b\right)$ are sufficient for the Radon-Nikodym derivatives of $P_{XY}$ with respect to $P_{X}P_{Y}$; and $P_{YX}$ with respect to $P_{Y}P_{X}$ to exist. Hence, it follows that for all sets $\mathcal{A}\leavevmode\nobreak\ \in\leavevmode\nobreak\ \mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}$,
@@
$P_{XY}\left(\mathcal{A}\right)$ $=$ $\iint_{\mathcal{A}_{y}}\mathrm{d}P_{X|Y=y}\left(x\right)\mathrm{d}P_{Y}\left(y\right)$ (43)
$=$ $\iint_{\mathcal{A}_{y}}\frac{\mathrm{d}P_{X|Y=y}}{\mathrm{d}P_{X}}\left(x\right)\mathrm{d}P_{X}\left(x\right)\mathrm{d}P_{Y}\left(y\right)$ (44)
$=$ $\iint\mathds{1}_{\mathcal{A}_{y}}(x)\frac{\mathrm{d}P_{X|Y=y}}{\mathrm{d}P_{X}}\left(x\right)\mathrm{d}P_{X}\left(x\right)\mathrm{d}P_{Y}\left(y\right)$ (45)
@@
$P_{XY}\left(\mathcal{A}\right)$ $=$ $\iint_{\mathcal{A}_{x}}\mathrm{d}P_{Y|X=x}\left(y\right)\mathrm{d}P_{X}\left(x\right)$ (48)
$=$ $\iint_{\mathcal{A}_{x}}\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)\mathrm{d}P_{Y}\left(y\right)\mathrm{d}P_{X}\left(x\right)$ (49)
$=$ $\iint\mathds{1}_{\mathcal{A}_{x}}(y)\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)\mathrm{d}P_{Y}\left(y\right)\mathrm{d}P_{X}\left(x\right)$ (50)
$=$ $\iint\mathds{1}_{\mathcal{A}_{y}}(x)\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)\mathrm{d}P_{X}\left(x\right)\mathrm{d}P_{Y}\left(y\right)$ (51)
@@
$P_{XY}\left(\mathcal{A}\right)=\int_{\widehat{\mathcal{A}}}\mathrm{d}P_{YX}\left(y,x\right)$ (54)
$=$ $\int_{\widehat{\mathcal{A}}}\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{Y}P_{X}\left(y,x\right)$
$=$ $\int\mathds{1}_{\widehat{\mathcal{A}}}(y,x)\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{Y}P_{X}\left(y,x\right)$ (55)
$=$ $\iint\mathds{1}_{\widehat{\mathcal{A}}_{x}}(y)\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{Y}(y)\mathrm{d}P_{X}(x)$ (56)
$=$ $\iint\mathds{1}_{\widehat{\mathcal{A}}_{y}}(x)\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{X}(x)\mathrm{d}P_{Y}(y)$ (57)
$=$ $\int\mathds{1}_{\widehat{\mathcal{A}}}(x,y)\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{X}P_{Y}(x,y)$ (58)
@@
where the set $\widehat{\mathcal{A}}$ is defined in (27). The equality in (54) is obtained by performing a change of measure using Theorem 2 under assumption $\left(a\right)$; and the equality in (57) follows by exchanging the order of integration [5, Theorem 2.6.6].
The proof is completed from Theorem 1 and by combining equations (42),  (47),  (53) and (59), which establish (40) and (41). ∎
Theorem 11 (Inverse Bayes-like Rule).
Consider the conditional probability measures $P_{Y|X}$ and $P_{X|Y}$; and the probability measures $P_{Y}$ and $P_{X}$ that satisfy (31) and (32); and the joint probability measures $P_{YX}$ and $P_{XY}$ in (26) and (33) respectively. Assume that: $\left(a\right)$ For all $x\in\mathcal{X}$${P_{Y}}\ll{P_{Y|X=x}}$; and $\left(b\right)$ For all $y\in\mathcal{Y}$${P_{X}}\ll{P_{X|Y=y}}$. Then,
$\frac{\mathrm{d}P_{X}P_{Y}}{\mathrm{d}P_{XY}}\left(x,y\right)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{XY}}\frac{\mathrm{d}P_{X}}{\mathrm{d}P_{X|Y=y}}\left(x\right)\frac{\mathrm{d}P_{Y}}{\mathrm{d}P_{Y|X=x}}\left(y\right)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{YX}}\frac{\mathrm{d}P_{Y}P_{X}}{\mathrm{d}P_{YX}}\left(y,x\right).$
@@
The following theorem introduces an identity involving the sum of a mutual information [13, 14] and a lautum information [33]. Such a result appeared first in [42, Corollary 13] as a byproduct of the properties of the generalization error of machine learning algorithms. In this work, a novel proof is presented in a general context exclusively using the folklore theorems on the RND described above.
Theorem 12.
Consider the conditional probability measure $P_{Y|X}$; the probability measures $P_{Y}$ and $P_{X}$ that satisfy (31) and (32); and the joint probability measure $P_{YX}$ in (26). Let $Q$ be some $\sigma$-finite measure in $\triangle\left(\mathcal{Y}\right)$ and assume that for all $x\in\mathcal{X}$$P_{Y}\ll P_{Y|X=x}\ll Q\ll P_{Y}$. Then,
$\mathrm{I}\left(P_{Y|X};P_{X}\right)+\mathrm{L}\left(P_{Y|X};P_{X}\right)$ (60)
$=$ $\iint\log\left(\displaystyle\frac{\mathrm{d}{P_{Y|X=x}}}{\mathrm{d}{Q}}(y)\right)\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x)$
-- $\iint\log\left(\displaystyle\frac{\mathrm{d}{P_{Y|X=x}}}{\mathrm{d}{Q}}(y)\right)\mathrm{d}P_{Y}(y)\mathrm{d}P_{X}(x),$
where $\mathrm{I}\left(P_{Y|X};P_{X}\right)$ and $\mathrm{L}\left(P_{Y|X};P_{X}\right)$ are respectively the mutual information and lautum information induced by the joint probability measure $P_{YX}$ in (26).
Proof:
From the definition of mutual information, it follows that
$\mathrm{I}\left(P_{Y|X};P_{X}\right)=\iint\log\left(\displaystyle\frac{\mathrm{d}{P_{Y|X=x}}}{\mathrm{d}{P_{Y}}}(y)\right)\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x)$ (61)
$=$ $\iint\log\left(\displaystyle\frac{\mathrm{d}{P_{Y|X=x}}}{\mathrm{d}{Q}}(y)\displaystyle\frac{\mathrm{d}{Q}}{\mathrm{d}{P_{Y}}}(y)\right)\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x)$
$=$ $\iint\log\left(\displaystyle\frac{\mathrm{d}{P_{Y|X=x}}}{\mathrm{d}{Q}}(y)\right)\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x)$
$+$ $\iint\log\left(\displaystyle\frac{\mathrm{d}{Q}}{\mathrm{d}{P_{Y}}}(y)\right)\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x),$ (63)
where the equality in (IV) follows from Theorem 4. The proof follows by re-writing the second term in (63). That is,
$\iint\log\left(\displaystyle\frac{\mathrm{d}{Q}}{\mathrm{d}{P_{Y}}}(y)\right)\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x)$ (64)
$=$ $\iint\left(\log\left(\displaystyle\frac{\mathrm{d}{Q}}{\mathrm{d}{P_{Y}}}(y)\right)\displaystyle\frac{\mathrm{d}{P_{Y|X=x}}}{\mathrm{d}{P_{Y}}}(y)\right)\mathrm{d}P_{Y}(y)\mathrm{d}P_{X}(x)$
$=$ $\int\log\left(\displaystyle\frac{\mathrm{d}{Q}}{\mathrm{d}{P_{Y}}}(y)\right)\left(\int\displaystyle\frac{\mathrm{d}{P_{Y|X=x}}}{\mathrm{d}{P_{Y}}}(y)\mathrm{d}P_{X}(x)\right)\mathrm{d}P_{Y}(y)$ (65)
$=$ $\int\log\left(\displaystyle\frac{\mathrm{d}{Q}}{\mathrm{d}{P_{Y}}}(y)\right)\mathrm{d}P_{Y}(y),$ (66)
where the equality in (64) follows from Theorem 2; the equality in (65) is a reorganization of terms in the integrals; and the equality in (66) follows from Theorem 9. From the definition of lautum information, it follows that
$\mathrm{L}\left(P_{Y|X};P_{X}\right)=\iint\log\left(\displaystyle\frac{\mathrm{d}{P_{Y}}}{\mathrm{d}{P_{Y|X=x}}}(y)\right)\mathrm{d}P_{Y}(y)\mathrm{d}P_{X}(x)$
$=$ $\iint\log\left(\displaystyle\frac{\mathrm{d}{P_{Y}}}{\mathrm{d}{Q}}(y)\displaystyle\frac{\mathrm{d}{Q}}{\mathrm{d}{P_{Y|X=x}}}(y)\right)\mathrm{d}P_{Y}(y)\mathrm{d}P_{X}(x)$
$=$ $\int\log\left(\displaystyle\frac{\mathrm{d}{P_{Y}}}{\mathrm{d}{Q}}(y)\right)\mathrm{d}P_{Y}(y)$
$-\iint\log\left(\displaystyle\frac{\mathrm{d}{P_{Y|X=x}}}{\mathrm{d}{Q}}(y)\right)\mathrm{d}P_{Y}(y)\mathrm{d}P_{X}(x),$ (68)
where (IV) follows from Theorem 4 and (68) is a direct consequence of the properties of the logarithmic function and Theorem 5. The proof is completed by plugging (66) in (63); and adding the resulting (63) to (68). ∎
The relevance of expressing the sum of a mutual information and a lautum information in terms of a free parameter (i.e., the measure $Q$ in (60)) lies in a hypothesis-testing interpretation for the term $\log\left(\displaystyle\frac{\mathrm{d}{P_{Y|X=x}}}{\mathrm{d}{Q}}(y)\right)$, which can be viewed as a log-likelihood ratio for distinguishing the hypothesis $Y\sim P_{Y|X=x}$ from $Y\sim Q$, given some data point $(x,y)$. Consequently, such a sum captures the variation of the expectation of such a log-likelihood ratio when the ground truth probability measure shifts from $P_{Y}P_{X}$ to $P_{XY}$. This interpretation has played a central role for establishing bridges between the generalization theory and hypothesis testing [42].